OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, 1) and (4, 13)? (1 points) Question 15 options: 1) y - 1 = -2(x - 4) 2) y - 1 = 2(x + 2) 3) y - 13 = -2(x + 2) 4) y - 13 = 2(x + 4)

OpenStudy (anonymous):

2)

myininaya (myininaya):

\[m=\frac{y-y_1}{x-x_1}\] so \[(x-x_1)m=y-y_1\] so what is m? \[m=\frac{13-1}{4-(-2)}=\frac{12}{6}=2\] y-1=2(x-(-2)) y-1=2(x+2)